For a continuous function f(x) =kx2e-x where x>0.Find i) k, ii) mean, iii) variance

1
Expert's answer
2021-05-04T13:05:31-0400

"f(x) =kx^2e^{-x}, \; x>0"


i) "\int f(x) dx =1"

"\int_0^\infty kx^2 e^{-x}dx = k \int_0^\infty x^2 e^{-x} dx =/ \text{integrating by parts}/ = 2k =1"

"k=1\/2"


ii) mean

"M[x] = \int_0^\infty xf(x) dx= k\int_0^\infty x^3 e^{-x} dx = / \text{integrating by parts}/ = 6k =3"

"M[x] = 3"


iii) variance

"D[x] = M[x^2] - (M[x])^2"

"M[x^2] = \int_0^\infty x^2f(x) dx= k\int_0^\infty x^4 e^{-x} dx = / \text{integrating by parts}/ = 24k =12"

"D[x] = 12 - (3)^2 = 12 - 9 = 3"